# NEDS-SLAM: A Novel Neural Explicit Dense Semantic SLAM Framework using 3D Gaussian Splatting

### v2

@@ III-B3 Vitrual Camera Pruning 3D Gaussians
The key aspects of GS-based SLAM are: 1) Distinguishing established high-quality areas from areas that need further optimization, and 2) Identifying and removing outlier points. The former resolves where to add Gaussians, and also plays a key role in camera tracking. Areas of low quality can severely affect the accuracy of pose tracking. The second key aspect resolves where to delete Gaussians. Outlier points will cause holes and defects during image rendering, and these flaws can also affect the accuracy of camera tracking.
@@
The distinction between well-optimized and areas with low quality is implemented through Eq. 5. This section discusses issues related to Gaussians pruning.
MonoGS [13] employs multi-view consistency to eliminate outliers. A Gaussian is deemed visible from a particular view if it is used in rasterization and the ray’s cumulative alpha value has not yet surpassed 0.5. MonoGS maintains a keyframe window, and if Gaussians inserted within the last three keyframes are not observed by at least three other frames in this window, they are considered outliers. Consequently, these Gaussians are removed during the optimization process.
@@
Figure 3: The concept of virtual view pruning for identifying outlier Gaussian points. We analyze only the GS points visible in the current ground-truth view (points $A$, $B$, $C$ in the figure). Point A is not visible from either of the two virtual viewpoints, thus identified as an outlier GS point, and its opacity is degraded during subsequent optimization. While the figure depicts two virtual viewpoints in a planar scenario, our approach creates four virtual cameras by rotating the camera pose from the focal point of each GT view frame along four directions: up, down, left, and right.
@@
Figure 4: Rendered virtual camera views on the ScanNet dataset. he middle images provide a zoomed-in illustration of the effectiveness of Virtual Camera Pruning, where ’vv’ denotes virtual camera view. Eliminating outlier Gaussians not only improves rendering quality but also reduces the storage footprint of the map representation.
@@
Inspired by MonoGS, the proposed NEDS-SLAM introduces a novel Virtual Multi-View Consistency Check approach, as depicted in Fig.3. The points $A$ and $B$ represent outlier GS points, while the GT view denotes the camera pose estimated within the RGBD stream. In the current keyframe, both $A$ and $B$ are visible. However, in the left virtual viewpoint, neither of these outlier points is visible, and in the right virtual viewpoint, only $B$ is visible while $A$ is not. The virtual camera operates alongside the real camera. If a GS point is invisible in all virtual views but visible in the real view, it is then considered an outlier.
@@
The virtual multi-view consistency check method takes advantage of the fast rendering capabilities of the Gaussian Splatting model, enabling the marking of GS points that significantly deviate from the object surface. In subsequent optimization processes, the involvement of outlier GS in the scene is diminished by degrading their opacity. Consistent with [10], GS with near-zero opacity or excessive radius are removed in the mapping process. FSGS [22] used a similar method to improve the quality of images synthesized from novel views. Specifically, FSGS uses the new viewpoint to guide the optimization of grown Gaussians towards a reasonable geometry. Meanwhile, the proposed method in this paper uses multiple synthesized images to eliminate inaccurate Gaussians. As illustrated in Figure. 4, we render virtual views and further optimize the GS distribution only for keyframes. The specific approach for generating virtual views is not fixed. Although Gaussian splatting enables extremely fast virtual view synthesis (nearly 300 FPS), introducing too many viewpoints can compromise the system’s real-time performance. We choose four virtual views along the up, down, left, and right directions, which achieves a desirable balance between effectiveness and efficiency.
@@ III-B4 Camera tracking
The camera tracking phase involves estimating the relative pose of the camera for each new frame, based on the already established map model. The camera pose for the new frame is initialized under the assumption of constant velocity, which includes both a constant linear and angular velocity, as Eq. 9. $T$ and $R$ respectively denote translational and rotational displacement.
$\left[{{T_{t+1}},{R_{t+1}}}\right]=\left[{{T_{t}},{R_{t}}}\right]+\left[{{T_{t}},{R_{t}}}\right]-\left[{{T_{t-1}},{R_{t-1}}}\right]$ (9)
@@
${L_{tracking}}=\left[{{\lambda_{c}}{L_{c}}+{\lambda_{d}}{L_{d}}+{\lambda_{s}}{L_{1}}\left({{S_{render}},{S_{head}}}\right)}\right]\cdot M$ (10)
$M$ in Eq. 10 is computed as Eq. 5. Artifacts and flaws such as holes and spurious effects caused by outlier gaussians significantly impact the precision of camera tracking. Subsequent experiments demonstrate that the incorporation of semantic loss improve the tracking accuracy. This improvement is attributed to the enriched understanding of the geometric information of objects, facilitated by the integration of semantic features.
Figure 5: The first column shows the RGB reconstruction results. The second column shows the ground truth semantic labels. The third column shows the semantic labels predicted directly on the current frame using M2F [18]. The fourth column shows the semantic reconstruction results using the SGS-SLAM [4] method based on Splatam [11]. The fifth column shows the reconstruction results of our proposed model.
Methods Depth L1[cm] $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ ATE RMSE[cm] $\downarrow$
@@
Table I: Quantitative comparison of map reconstruction and localization accuracy between our proposed NEDS-SLAM and other NeRF-based dense SLAM methods, averaged over 8 scenes from the Replica dataset.
@@
NICE-SLAM 12.00 10.90 13.40 6.20 10.63
Vox-Fusion 68.84 27.28 23.30 9.41 32.21
Point-SLAM 10.24 22.16 14.77 9.54 14.18
SplaTAM 12.56 11.09 11.07 7.46 10.54
@@
Table II: We quantitatively compare our proposed NEDS-SLAM method against other radiance field-based SLAM on the ScanNet dataset, reporting the tracking metric of camera pose RMSE measured in centimeters.
@@ IV-A Experimental Setup
Metrics. We employ several metrics to evaluate the reconstruction quality in our study. These include PSNR, Depth-L1 (on 2D depth maps), SSIM, and LPIPS. Additionally, we assess the accuracy of camera pose estimation using the average absolute trajectory error (ATE RMSE). To evaluate the performance of semantic segmentation, we calculate the mIoU (mean Intersection over Union) score.
Baselines. We compare the tracking and mapping with state-of-the-art methods iMAP, NICE-SLAM, Co-SLAM, ESLAM, and SplaTAM. For semantic segmentation accuracy, we compare with NIDS-SLAM, DNS-SLAM, and SNI-SLAM.
Implementation Details. We conducted experiments using a single NVIDIA 4090 GPU, validating on the REPLICA dataset with the mapping iteration set to 40, tracking iteration set to 60, and SCFF iteration set to 50. After obtaining 384 feature channels through the DINO model, we derived 64-dimensional fused features by applying 2D convolutions separately to the Spatial Features. Finally, we obtained three-dimensional features by passing them through an encoder and embedding them into the GS parameters. We use a learning rate of 0.005 and 0.001 respectively for all learnable parameters on Replica and ScanNet datasets. For camera poses, we only employ a learning rate of 0.0005 in tracking.
Metrics Depth L1[cm] $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ATE RMSE[cm] $\downarrow$ tracking/frame[s]$\downarrow$ mapping/frame[s] $\downarrow$
Room0 0.31 35.23 0.979 0.082 0.368 0.86 1.52
Room1 0.45 34.86 0.862 0.075 0.403 0.59 1.47
Room2 0.42 35.16 0.983 0.071 0.326 0.67 1.68
Office0 0.31 37.53 0.981 0.091 0.354 0.53 1.71
Office1 0.27 39.71 0.979 0.087 0.284 0.65 1.27
Office2 0.48 32.68 0.973 0.079 0.302 0.60 1.63
Office3 0.58 31.07 0.968 0.103 0.318 0.59 1.47
Office4 0.61 31.82 0.973 0.113 0.473 0.72 1.52
AVG 0.47 34.76 0.962 0.088 0.354 0.65 1.53
Table III: Detailed test results of NEDS-SLAM across 8 scenes from the Replica dataset.
[…]
base model with M2F head 33.57 0.372 26.52
[…]
with SL 32.52 0.368 30.28(+14%)
with FF 33.57 0.371 41.31(+50%)
[…]
Table IV: An ablation study of the NEDS-SLAM, where ’VV’ represents the virtual view pruning method, ’FF’ represents the spatial-consistent semantic fusion method, and ’SL’ represents the utilizing of semantic loss during training.
@@ IV-B Experiment result
Quantitative measures of reconstruction quality using the Replica dataset are presented in Table I. Our method demonstrates competitive performance when compared to other approaches.
The NEDS-SLAM, built upon the foundation of 3DGS, achieves accurate camera localization and semantic reconstruction simultaneously. Table. IV provides a comparison between our method and other neural Implicit approaches in terms of semantic reconstruction performance. Due to the precise representation of object edges offered by the Gaussian radiance field, methods based on neural explicit approaches bring about significant improvements in semantic reconstruction.
Semantic SLAM methods in real-world scenarios often employ a pretrained model to perform semantic segmentation on the RGB-D frames. This semantic information is then utilized in conjunction with geometric information to reconstruct the semantic scene. In an ideal scenario, a perfect semantic segmentation model would exhibit excellent spatio consistency. This means that the same object would be assigned accurate and consistent semantic features across different viewpoints and time instances.
However, the reality often falls short of perfection, and the limitations of semantic SLAM become apparent in the following aspects:
@@
NIDS-SLAM 82.37 82.45 84.08 85.94
DNS-SLAM 84.77 88.32 84.90 84.66
SNI-SLAM 87.41 88.42 87.43 87.63
@@
Table V: A quantitative comparison of our method with existing semantic NeRF-based SLAM approaches on the Replica dataset,For consistency and fair comparison with other methods, we utilize the ground-truth semantic labels from the replica dataset instead of the predicted semantic labels from pre-trained models
  • Inaccurate estimation across frames: Semantic segmentation models may struggle to accurately estimate semantic information across consecutive frames, leading to inconsistencies in the semantic understanding of the scene.
  • Trade-off between performance and inference time: Higher-performing models often require more computational resources and inference time, which can hinder the real-time nature of SLAM systems. The size and complexity of the semantic segmentation model must be carefully balanced to ensure efficient and timely processing.
When testing the dinov2_vit14 model on the replica room0 scene, as shown in Figure. 5, there are noticeable inconsistencies in the predictions for the floor and chairs. This affects the semantic reconstruction quality, but does not impact the reconstruction of the RGB channels. As shown in Figure. 5, NEDS-SLAM effectively filters out the negative impact of spatial semantic inconsistencies, generating robust semantic estimates and providing more accurate semantic reconstruction.
@@ IV-C Ablation Study
Ablation experiments were designed for the Feature Fusion module and Pruning GS module to test the effectiveness of NEDS-SLAM for semantic reconstruction. As shown in Table. IV, following SGS-SLAM’s approach, we directly incorporated semantic parameters into the GS by calling a pre-trained M2F segmentation model on each RGB frame. The segmented pixel class labels were color-coded and embedded into the GS parameters, resulting in the reconstruction effect displayed in the fourth column of Figure. 5, corresponding to the first row of Table. IV. For the Replica Room0 dataset, the M2F model achieved a semantic segmentation mIoU of 52.4. Employing this method for semantic SLAM resulted in an average mIoU of 26.52, which was used as the baseline approach. After incorporating the virtual view pruning method, outlier GS points that affected reconstruction quality were effectively removed. This led to a 5% increase in PSNR, slightly improved camera pose tracking accuracy, while maintaining the quality of semantic reconstruction. Building upon the baseline, the incorporation of a semantic loss along with depth and RGB losses, weighted at 0.5, 1.0, and 1.0 respectively, resulted in a 14% improvement in semantic reconstruction performance. The addition of the spatial-consistent semantic fusion method, which combined the semantic features from the DinoV2 ViT-14 model with the appearance features from the DepthAnything model, led to nearly a 50% improvement in semantic reconstruction performance. A lightweight encoder-decoder compressed the high-dimensional fused features into a low-dimensional representation, which was then embedded into the GS parameters.
As can be seen in Figure. 6, the semantic features calculated by the M2F model were inconsistent (such as the partitions and books on the table). After processing with the SCFF module, the inconsistencies were resolved and NEDS-SLAM output a more complete semantic reconstruction.
The baseline method suffered from significant semantic noise, resulting in a low semantic reconstruction score. In contrast, the proposed spatial-consistent semantic fusion method effectively addressed the issue of semantic inconsistency, leading to a remarkable improvement in reconstruction performance.

### v3

@@ III-B3 Vitrual Camera View Pruning 3D Gaussians (VCVP)
The key aspects of GS-based SLAM are: 1) Distinguishing well-established areas from areas requiring further optimization, and 2) Identifying and removing outlier points. The former resolves where to add gaussians, and also plays a key role in camera tracking. Areas of low quality can severely affect the accuracy of pose tracking. The second key aspect resolves where to delete gaussians. Outlier points will cause holes and defects during image rendering, and these flaws can also affect the accuracy of camera tracking.
@@
The distinction between well-optimized and areas with low quality is implemented through Eq 6. This section discusses issues related to gaussians pruning.
Multi-view consistency constraints have been proven effective in identifying geometrically unstable gaussians. Previous methods [4] check whether gaussians inserted within the latest three frames of a keyframe window are recorded by other keyframes, thereby determining outlier gaussians. This method improves mapping accuracy by using collaborative constraints among multiple keyframes, but it increases computational costs and reduces real-time performance. Drastic viewpoint changes during SLAM cause significant overlap variations between keyframes, leading to errors in outlier detection.
@@
Figure 2: The concept of virtual view pruning for identifying outlier gaussians. We analyze only the gaussians visible in the current ground-truth view (points $A$, $B$, $C$ in the figure). Point A is not visible from either of the two virtual views, thus identified as an outlier gaussians, and its opacity is degraded during subsequent optimization. While the figure depicts two virtual views in a planar scenario, our approach creates four virtual cameras by rotating the camera pose from the focal point of each GT view frame along four directions: up, down, left, and right.
@@
Figure 3: Rendered virtual camera views on the ScanNet dataset. The middle images provide a zoomed-in illustration of the effectiveness of Virtual Camera Pruning, where ’vcvp’ denotes virtual camera view. Eliminating outlier gaussians not only improves rendering quality but also reduces the storage footprint of the map representation.
In contrast to this method, the VCVP method proposed in this paper does not perform comparison between keyframes. Instead, it compares the viewpoint between a real camera frame and the corresponding virtual accompanying camera frame, as depicted in Fig 2. The virtual cameras ($VC$) are created by rotating the real camera ±$\pm\theta$ around the focal point: $VC_{1}$ and $VC_{2}$ by rotating on the horizontal plane ($xz$ plane), and $VC_{3}$ and $VC_{4}$ by rotating on the vertical plane ($yz$ plane).
@@
The points $A$ and $B$ represent outlier gaussians, while the GT view denotes the camera pose estimated within the RGBD stream. In the current keyframe, both $A$ and $B$ are visible. However, in the $VC_{1}$, neither of these outlier points is visible, and in the $VC_{2}$, only $B$ is visible while $A$ is not. The virtual camera operates alongside the real camera. If a gaussians is invisible in all virtual views but visible in the real view, it is then considered an outlier.
@@
The virtual multi-view consistency check method takes advantage of the fast rendering capabilities of the Gaussian Splatting, enabling the marking of gaussians that significantly deviate from the object surface. The VCVP method eliminates the dependence on historical keyframes, allowing it to remain unaffected by drastic changes in camera views. This enables the identification of single-view outlier gaussians. In subsequent optimization processes, the involvement of outlier gaussians in the scene is diminished by degrading their opacity. Consistent with [14], gaussians with near-zero opacity or excessive radius are removed in the mapping process. As illustrated in Fig 3, we render virtual views and further optimize the 3D gaussians parameters only for keyframes. The specific approach for generating virtual views is not fixed. Although Gaussian splatting enables extremely fast virtual view synthesis (nearly 300 FPS), introducing too many viewpoints can compromise the system’s real-time performance. We conducted detailed tests in Section IV-C to evaluate how the generation and function of the virtual camera impact the performance of the SLAM system. We choose four virtual views along the up, down, left, and right directions, which achieves a desirable balance between effectiveness and efficiency.
@@ III-B4 Camera tracking
The camera tracking phase involves estimating the relative pose of the camera for each new frame, based on the already established map model. The camera pose for the new frame is initialized under the assumption of constant velocity, which includes both a constant linear and angular velocity.
@@
$L_{\text{tracking}}=\left(\lambda_{c}L_{c}+\lambda_{d}L_{d}+\lambda_{s}L_{\text{CE}}\left(S_{\text{render}},S_{\text{head}}\right)\right)\cdot M$ (10)
$M$ in Eq 10 is computed as Eq 6. Artifacts and flaws such as holes and spurious effects caused by outlier gaussians significantly impact the precision of camera tracking. Subsequent experiments demonstrate that the incorporation of semantic loss improve the tracking accuracy. This improvement is attributed to the enriched understanding of the geometric information of objects, facilitated by the integration of semantic features.
Figure 4: The first row shows the RGB reconstruction results. The second row shows the semantic labels predicted directly on the current frame using M2F[24]. The third row shows the semantic reconstruction results using the SGS-SLAM[6] method based on SplaTAM[5]. The fourth row shows the reconstruction results of our proposed model.
Table I: Comparison experiments with other methods on map reconstruction and localization accuracy
[…]
@@
Table II: Comparison experiment on the ATE RMSE metric
@@
NICE-SLAM[3] 12.00 10.90 13.40 6.20 10.63
Vox-Fusion[25] 68.84 27.28 23.30 9.41 32.21
Point-SLAM[28] 10.24 22.16 14.77 9.54 14.18
SplaTAM[5] 12.56 11.09 11.07 7.46 10.54
@@ IV-A Experimental Setup
Metrics. We employ several metrics to evaluate the reconstruction quality in our study. These include the peak signal-to-noise ratio (PSNR), Depth-L1 (on 2D depth maps), Structural Similarity (SSIM[31]), and Learned Perceptual Image Patch Similarity (LPIPS[32]). Additionally, we assess the accuracy of camera pose estimation using the average absolute trajectory error (ATE RMSE [33]). To evaluate the performance of semantic segmentation, we calculate the mIoU (mean Intersection over Union) score.
Baselines. We compare the tracking and mapping with state-of-the-art methods NICE-SLAM[3], Co-SLAM[26], ESLAM[27], and SplaTAM[5]. For semantic segmentation accuracy, we compare with NIDS-SLAM[12], DNS-SLAM[11], and SNI-SLAM[10].
Implementation Details. We conducted experiments using a single NVIDIA RTX 4090 and an Intel Xeon Platinum 8358P, validating on the REPLICA dataset with the mapping iteration set to 40, tracking iteration set to 60, and SCFF iteration set to 50. After obtaining 384 feature channels through the DINO model, we derived 64-dimensional fused features by applying 2D convolutions separately to the Spatial Features. Finally, we obtained three-dimensional features by passing them through an encoder and embedding them into the gaussians parameters. We use a learning rate of 0.005 and 0.001 respectively for all learnable parameters on Replica and ScanNet datasets. For camera poses, we only employ a learning rate of 0.0005 in tracking.
Figure 5: The comparison validated the effectiveness of VCVP method. Showing that NEDS-SLAM achieves better reconstruction results in details compared to Splatam.
@@ IV-B Experiment result
Quantitative measures of reconstruction quality using the Replica dataset are presented in Table I. The experiments on the ScanNet dataset can be found in Table II. The data shows that our method achieves the highest camera pose tracking accuracy. Our method demonstrates competitive performance when compared to other approaches. As shown in Fig 5, due to the VCVP method removing geometrically unstable gaussians, our approach is able to preserve more details.
The NEDS-SLAM, built upon the foundation of 3DGS, achieves accurate camera localization and semantic reconstruction simultaneously. Table III provides a comparison between our method and other neural Implicit approaches in terms of semantic reconstruction performance.
Due to the precise representation of object edges offered by the 3DGS, our methods bring about significant improvements in semantic reconstruction. Other methods have not considered the issue of spatially inconsistent semantic estimation by pre-trained semantic segmentation models on consecutive RGBD frame inputs. Therefore, for the sake of fair performance comparison in Table III, we used the ground truth per-pixel semantic class labels as input. More detailed experiments on the SCFF module are conducted in Table V in Section IV-C.
Table III: Comparison experiment on the mIoU metric
@@
NIDS-SLAM[12] 82.37 82.45 84.08 85.94
DNS-SLAM[11] 84.77 88.32 84.90 84.66
SNI-SLAM[10] 87.41 88.42 87.43 87.63
@@
When testing the Mask2Former model on the replica room0 scene, as shown in Fig 4, there are noticeable inconsistencies in the predictions for the floor and chairs. This affects the semantic reconstruction quality. As shown in Fig 4, NEDS-SLAM effectively filters out the negative impact of spatial semantic inconsistencies, generating robust semantic estimates and providing more accurate semantic reconstruction.
@@ IV-C Ablation Study
Effectiveness of VCVP Module.
The VCVP method involves two subproblems: (1) determining the number of virtual camera views to generate and how to generate them, and (2) deciding on which frame or frames to perform VCVP operation. The solutions to these subproblems will impact the computational costs of the VCVP modules. In Table IV, the data in the third and fourth rows labeled ’A/B/C’ indicates that we used three configurations for the calculations. Configuration A and B represent generating two virtual views in the horizontal and vertical directions, respectively. Configuration C represents generating four virtual views simultaneously in both horizontal and vertical directions.
The VCVP module significantly enhances scene modeling accuracy and camera pose tracking precision. Increasing the number of virtual camera views and their usage within the keyframes window can achieve the best camera localization accuracy, but this also increases computational overhead. In our most extreme test case, VCVP detection was performed on 10 keyframes during each mapping iteration, with four virtual camera viewpoints rendered for each detection, resulting in nearly a 50% improvement in pose accuracy
We conducted another experiment comparing our method to the density control method (denote as DC method) from the original 3DGS paper, as in Table.VII. In experiments on three scenes from the TUM RGBD dataset, the DC method achieved ATE RMSE values of 3.62, 1.41, and 6.63, which are higher than those of GS-SLAM, which also uses partial DC operations (3.3, 1.3, and 6.6 respectively). Our VCVP method demonstrated even higher performance.
Settings
ATE
RMSE $\downarrow$
AVG
SSIM $\uparrow$
Scene
Embedding $\downarrow$
Mapping
/iteration $\downarrow$
Base1 0.42 0.90 100.16 MB 14 ms
Base + VCVP_w5* 0.30/0.34/0.28 0.91/0.93/0.96 90.45 MB 16/16/20 ms
Base + VCVP_w10 0.26/0.27/0.22 0.92/0.92/0.97 88.93 MB 18/18/26 ms
Base + RCVP2 0.36 0.95 95.27 MB 20 ms
[…]
Co-SLAM[26] 0.97 0.91 - 13 ms
NICE-SLAM[3] 0.99 0.69 48.48 MB 66 ms
  • 1
    Base refers to the configuration without SCFF, without lightweight encoder, and without VCVP, implementing only the 3DGS dense SLAM functionality.
    VCVP_w5(10) indicates selecting 5(10) frames from the current keyframes window for VCVP operations.
  • 2
    RCVP involves using real camera views for consistency checks and removing outlier gaussians.
Table IV: Ablation experiments on the VCVP module conducted in replica room0.
Settings mIoU  $\uparrow$
Mapping
/iteration $\downarrow$
Scene
Embedding $\downarrow$
Base_S 26.52% 28 ms 123.08 MB
[…]
Base_S + SCFF_wo_SFE
30.24% 86 ms 405.64 MB
[…]
Base_S + SCFF_w_SFE
42.18% 86 ms 410.38 MB
[…]
Base_S + SCFF_w_SFE + encoder
40.81% 35 ms 141.93 MB
Table V: Ablation Study of the SCFF Module on ScanNet dataset.
Table VI: Runtime performance comparison of NEDS-SLAM on two different hardware platforms.
Hardware Settings # replica room0 # TUM RGBD fr1/desk
Tracking/it $\downarrow$
Mapping/it $\downarrow$
Tracking/it $\downarrow$
Mapping/it $\downarrow$
Platform A 28 ms 35 ms 26 ms 34 ms
Platform B 42 ms 76 ms 42 ms 75 ms
Table VII: Comparison of the VCVP module with the original density control method on TUM-RGBD dataset.
DATASETS with VCVP Original density control method as in [14]
[…]
Fr1/desk1 3.30 0.91 3.62 0.93
Fr2/xyz 1.13 0.95 1.41 0.95
Fr3/off 4.94 0.90 6.63 0.92
Table VIII: Verification of the effectiveness of the SCFF module
Model Settings M2F[24] M2F+SCFF MRCNN[34] MRCNN+SCFF
AVG mIoU 25.89% 36.25% 24.34% 34.07%
Effectiveness of SCFF Module.
Following SGS-SLAM, we directly incorporated semantic parameters into the 3D gaussians by calling a pre-trained M2F segmentation model [24] on each RGB frame. As shown in the third row in Fig 4, corresponding to the Base_S settings in Table V. SCFF_wo_SFE represents configurations includes the SCFF module, but does not use SFE. For the ScanNet scene0000 dataset, the M2F model achieved a semantic segmentation mIoU of 52.4. Using M2F for segmentation head gave an average mIoU of 26.52, serving as the baseline.
As can be seen in Fig 6, the semantic features calculated by the M2F model were inconsistent (such as the partitions and books on the table). After processing with the SCFF module, the inconsistencies were resolved and NEDS-SLAM output a more complete semantic reconstruction. The SCFF module filters out unstable semantic estimations between frames, resulting in more accurate semantic reconstruction. Our designed SCFF features a lightweight network structure, which does not significantly increase inference time. In fact, the time consumption in the SLAM process (Mapping/Iteration in the table) mainly arises from optimizing a large number of 3D gaussians. Therefore, our specially designed encoder compresses the semantic features and embeds them into the gaussian parameters, reducing the number of parameters and thereby increasing the mapping speed.
Runtime Comparison.
As shown in the last column of Table VI, our lightweight configuration of NEDS-SLAM achieves faster mapping speeds than SplaTAM while maintaining more accurate camera pose tracking precision. With higher configurations, NEDS-SLAM offers better performance, though the computation speed decreases. We ran NEDS-SLAM on different hardware platforms and datasets. Platform A is as in section IV-A. Platform B consists of an Intel i9-13900K and a single NVIDIA RTX 4060Ti. The results show that both hardware platforms achieve similar camera pose tracking accuracy and reconstruction accuracy with NEDS-SLAM, but the model takes more time to run on Platform B compared to Platform A.